OpenStudy (anonymous):

4 pencils have the same value of 3 pens. One pen is 1/3 the cost of stapler. Two staplers are worth 3 packets of notecards. How many pencils are worth one packet of notecards. Express your answer as a common fraction.

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

4p = 3p p = 1/3s 2s = 3n

OpenStudy (anonymous):

?

OpenStudy (igreen):

Lol

OpenStudy (igreen):

3 * 1/3 = ?

OpenStudy (anonymous):

8/3

OpenStudy (igreen):

Check again

OpenStudy (anonymous):

1

OpenStudy (anonymous):

@iGreen is it 1

OpenStudy (igreen):

Yes, it is.

OpenStudy (igreen):

So 3 pens are equal to 1 stapler.

OpenStudy (anonymous):

okay so that is the answer

OpenStudy (igreen):

And 4 pencils are equal to 3 pens. So 4 pencils are equal to 1 stapler.

OpenStudy (igreen):

2 staplers are equal to 3 notebook packets.

OpenStudy (igreen):

2 * 4 = ?

OpenStudy (anonymous):

8

OpenStudy (igreen):

Yes, so 8 pencils are equal to 2 staplers. So 8 pencils are also equal to 3 notebook packets.

OpenStudy (anonymous):

so 8/3

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

thx :)